# SOFIA/HAWC+ FAR-INFRARED POLARIMETRIC LARGE-AREA CMZ EXPLORATION (FIREPLACE) SURVEY III: FULL SURVEY DATA SET

### v2

@@ 5.3 Clouds Surrounding Sgr A
The core of the 20 km s cloud is comprised of a set of filamentary structures in the dust continuum that appear to merge at =359.867, =-0.083. At this location we observe low fractional polarization values (1%), suggesting that depolarization in this region may be present. This depolarization could be caused by the overlapping of field directions along our line-of-sight from these multiple filamentary structures. Alternative explanations for depolarization in this region could be caused by loss of grain alignment in the densest regions of the cloud or multiple field components within the size of our beam. Depolarization in these denser regions has been observed in several of CMZ clouds including Sgr B2 and Dust Ridge cloud E/F (FIREPLACE I). Previous studies using combined SMA and VLA observations find that the star formation of the 20 km s cloud is likely in an early evolutionary stage (Lu et al., 2015). Sgr B2 and the Dust Ridge cloud E/F, which also show depolarization in their cores show signs of early and on-going star formation - indicating that the 20 km s cloud may be at a similar evolutionary state.
@@
Sgr C is located 0.5 in projection from the 20 km s cloud and is on the Western periphery of the CMZ as shown in Figure 15. The cloud manifests a mix of diffuse and compact emission coupled with a broad range of emissivities as shown in Figure 20. This cloud was recently observed polarimetrically using the JCMT (Lu et al., 2023) and the CNM mode observations made using SOFIA/HAWC+ that are shown in Figure 6. Prior to these observations Sgr C had not been studied polarimetrically at a resolution less than 1′. The FIREPLACE observations of Sgr C yield 100 magnetic field pseudovectors within the periphery of the cloud. These observations largely agree with the orientations derived from the SOFIA/HAWC+ CNM observations of this cloud as seen in Figure 6.
The emission from Sgr C obtained by Herschel is generally diffuse, but there is a region of compact, high intensity emission at = 359.433, = -0.1. There is a ridge or tail of emission emanating from this compact emission towards the Southeast, which we hereafter refer to as the “head-tail structure” (labeled in Figure 20). The magnetic field derived from our 214 µm observations follows the morphology of the head-tail structure and is largely consistent with the magnetic field orientation derived from the 850 µm JCMT observations of Lu et al. (2023). Lu et al. (2023) argue that the magnetic field orientation in the head-tail structure could result from interaction of the Sgr C molecular cloud with a nearby prominent H II region, the extent of which has been marked with an orange circle in Figure 20 (Hankins et al., 2020; Lang et al., 2010). Evidence of this interaction between Sgr C and the H II region has previously been presented in (e.g. Lu et al., 2019). The similar magnetic field orientations derived from FIREPLACE bolster the argument that winds from the prominent H II region create molecular pillars leading to a compressed magnetic field that follows the orientation of the head-tail structure. Magnetohydrodynamic simulations of expanding H II regions reveal similar molecular and magnetic field morphologies to what is seen in the head-tail structure of Sgr C (Arthur et al., 2011; Mackey & Lim, 2011).
@@
The magnetic field revealed by our SOFIA/HAWC+ observations Southeast of the head-tail structure at = 359.467, = -0.15 is largely oriented perpendicular to the Galactic plane. The JCMT pseudovectors in this Southern portion of Sgr C generally do not agree with those obtained in the FIREPLACE survey and are more spatially varying. To help understand the difference in polarization angle between the FIREPLACE and JCMT observations, we note that the significance of the JCMT polarized intensity in this region is reported as , whereas the significance of the JCMT polarized intensity in the head-tail structure in Sgr C is (Lu et al., 2023). is equivalent to a 10 uncertainty in the polarization angle, and the JCMT orientations disagree with those obtained from FIREPLACE by 70-90 in this Southern region.
To the Southwest of the head-tail structure our SOFIA/HAWC+ observations reveal a curved magnetic field structure at = 359.417, = -0.167. There are also JCMT pseudovectors at this location which are again highly spatial variable. The curvature in this region is similar to the field structures seen for the Brick and the TLP clouds that were attributed to out-flowing winds. Furthermore, a compact source from the ATLASGAL survey (AGAL359.474-0.152; Contreras et al., 2013), whose location is marked with a red ellipse in Figure 20, has an integrated flux of 168 Jy and a total gas mass of 5 10 M (Kendrew et al., 2013). These properties make AGAL359.474-0.152 favorable for forming massive stars (Lada et al., 2012).
Therefore, out-flowing winds from star formation in Sgr C, likely originating from AGAL359.474-0.152 is a reasonable origin for the structure of the magnetic field in this region of Sgr C, especially in light of the ample evidence of star formation in Sgr C (e.g. Lu et al., 2019; Yusef-Zadeh et al., 2009). However, there are alternative mechanisms that could result in the magnetic field distribution seen in this region, such as shear from orbital motion of the structure.
As with the previously discussed clouds, the fractional polarizations derived from our 214 µm SOFIA/HAWC+ observations increase towards the periphery of Sgr C in lower emission regions. In the region of diffuse emission to the South of the head-tail structure fractional polarizations range from 4 – 5%. In the head-tail structure, conversely, the fractional polarizations are 1%, similar to the peak emission regions of other clouds.
@@ 6 CONCLUSIONS
In this work the full FIREPLACE survey covering the entire CMZ polarimetrically at 214 µm is presented. First, the reduction method performed on the observations obtained in 2022 is detailed, comprising the second FIREPLACE data release (DR2) and observing the region of the CMZ extending from the Brick to Sgr C (corresponding to a 1 0.75 region of the sky). The reduced DR2 observations are then combined with the DR1 observations presented in FIREPLACE I to obtain a map of the entire CMZ polarimetrically at 214 µm (a 1.5 0.75 region of the sky). CMZ-wide trends of the magnetic field are then inferred and analyzed, with the fields local to a set of prominent molecular clouds inspected in more detail. The following summarizes the results from the inspection of the full FIREPLACE survey:
@@
The sensitivity of our HAWC+ observations makes it possible to derive the polarization angle and magnetic field orientations in fainter molecular structures in the CMZ at a resolution of 19.6″ than ever before. The distribution of 65,000 Nyquist-sampled magnetic field pseudovectors from the full FIREPLACE survey is shown in Figure 10.
@@
A bimodal distribution in the CMZ-wide magnetic field orientations is identified from the full FIREPLACE survey. This bimodal distribution shows enhancements in directions parallel (approximately -18) and perpendicular (approximately -86) to the Galactic plane. This is a similar bimodal distribution to the one recovered in FIREPLACE I for the Sgr B2 region. The component that is parallel to the Galactic plane could either indicate a toroidal magnetic field component in the CMZ (Guan et al., 2021; Mangilli et al., 2019) or a component external to the CMZ. The bimodal component that is perpendicular to the Galactic plane could be related to the poloidal magnetic field distribution in the GC (e.g. Morris, 2006), possibly revealing connections between the magnetic fields traced by infrared dust polarization and radio synchrotron emission.
@@
The distributions of the magnetic fields of a set of prominent molecular clouds in the CMZ are also analyzed: the Brick, Three Little Pigs (TLP) cloud complex, 50 km s cloud, CND, CO 0.02-0.02, 20 km s cloud, and Sgr C. The magnetic field derived from FIREPLACE generally traces the morphology of the molecular clouds. In addition, there is evidence that expanding shock fronts from SNRs, advancing ionization fronts from H II regions, and tidal forces from orbital motion might cause compression and shearing leading to the organization of the magnetic fields in the Brick, TLP cloud complex, 50 km s cloud, and Sgr C. However, not all of the prominent clouds studied in this work exhibit signatures of compression or winds (the 20 km s cloud, for example). These results nonetheless indicate that the compression from winds and shear from tidal forces are important mechanisms for ordering the magnetic field distributions within CMZ clouds.
@@
The FIREPLACE observations presented in this work have the potential to enhance our understanding of the role magnetic fields play in the star formation process in the complex CMZ. Furthermore, the magnetic field distribution derived from these observations will improve our ability to develop a unified picture of the CMZ-wide magnetic field throughout the GC region.
@@ Appendix A.1 Guide to Individual Fields and CAL Files
We tabulate the fields and calibration files used for the DR2 FIREPLACE data set in Table A.2.1. The files in Table A.2.1 contain calibrated data with pointing corrections applied. Field numbers refer to the locations in Figure 2. “CCW” and “CW” refer to the rotation direction of the major scan direction relative to perpendicular to the Galactic North (30 as discussed in Section 2.1). For completeness, files from Flight 890 are included in this table, though as discussed in the main paper these files are not used in the DR2 reduction. Horizontal lines are included to separate observations from different flights.
Table A.2.2 tabulates the FIREPLACE DR1 files that are combined with the DR2 files shown in Table A.2.1 to produce the full FIREPLACE data release covering the entire CMZ. The format of Table A.2.2 is identical to that of Table A.2.1 except that orientation (“CCW” or “CW”) is not specified since the orientation of the DR1 observations was more variant, as discussed in Section 2.1.
@@ Appendix A.2 Histograms of the 2022 Polarization Pseudovectors
To do so we sort the Nyquist-sampled polarization angles shown in Figure 3 into 10 bins. The histogram representation of the binned polarization angles for the DR2 observations is shown in Figure A.1. As with the histogram of the full FIREPLACE release, an angle of 0 indicates a magnetic field that is aligned with the Galactic plane, with an angle of 90 indicating a magnetic field that is oriented perpendicular to the Galactic plane.

### v3

@@ 5.3 Clouds Surrounding Sgr A∗
The core of the 20 km s-1 cloud is comprised of a set of filamentary structures in the dust continuum that appear to coalesce at the brightest 250 µm clump in the cloud, located at =359.867∘, =-0.083∘. At this location we observe low fractional polarization values (1%), suggesting that depolarization in this region may be present. This depolarization could be caused by the overlapping of field orientations along our line-of-sight from these multiple filamentary structures. Alternative explanations for depolarization in this region could be caused by loss of grain alignment in the densest regions of the cloud or multiple field components within the size of our beam. Depolarization in these denser regions has been observed in several of CMZ clouds including Sgr B2 and Dust Ridge cloud E/F (FIREPLACE I). Previous studies using combined SMA and VLA observations find that the star formation of the 20 km s-1 cloud is likely in an early evolutionary stage (Lu et al., 2015). Sgr B2 and the Dust Ridge cloud E/F, which also show depolarization in their cores show signs of early and on-going star formation - indicating that the 20 km s-1 cloud may be at a similar evolutionary stage.
@@
Sgr C is located 0.5∘ in projection from the 20 km s-1 cloud and is on the Western periphery of the CMZ as shown in Figure 16. The cloud manifests a mix of diffuse and compact emission coupled with a broad range of emissivities as shown in Figure 21. This cloud was recently observed polarimetrically using the JCMT (Lu et al., 2024) and the CNM mode observations made using SOFIA/HAWC+ that are shown in Figure 6. Prior to these observations Sgr C had not been studied polarimetrically at a resolution less than 1′. The FIREPLACE observations of Sgr C yield 100 magnetic field pseudovectors within the periphery of the cloud. These observations largely agree with the orientations derived from the SOFIA/HAWC+ CNM observations of this cloud as seen in Figure 6.
The emission from Sgr C obtained by Herschel is generally diffuse, but there is a region of compact, high intensity emission at = 359.433∘, = -0.1∘. There is a ridge or tail of emission emanating from this compact emission towards the Southeast, which we hereafter refer to as the “head-tail structure” (labeled in Figure 21). The magnetic field derived from our 214 µm observations follows the morphology of the head-tail structure and is largely consistent with the magnetic field orientation derived from the 850 µm JCMT observations of Lu et al. (2024). Lu et al. (2024) argue that the magnetic field orientation in the head-tail structure could result from interaction of the Sgr C molecular cloud with a nearby prominent H II region, the extent of which has been marked with an orange circle in Figure 21 (Lang et al., 2010; Hankins et al., 2020). Evidence of this interaction between Sgr C and the H II region has previously been presented in e.g. Lu et al. (2019). The similar magnetic field orientations derived from FIREPLACE bolster the argument that winds from the prominent H II region create molecular pillars leading to a compressed magnetic field that follows the orientation of the head-tail structure. Magnetohydrodynamic simulations of expanding H II regions reveal similar molecular and magnetic field morphologies to what is seen in the head-tail structure of Sgr C (Arthur et al., 2011; Mackey & Lim, 2011).
@@
The magnetic field revealed by our SOFIA/HAWC+ observations Southeast of the head-tail structure at = 359.467∘, = -0.15∘ is largely oriented perpendicular to the Galactic plane. The JCMT pseudovectors in this Southern portion of Sgr C generally do not agree with those obtained in the FIREPLACE survey and are more spatially varying. To help understand the difference in polarization angle between the FIREPLACE and JCMT observations, we note that the significance of the JCMT polarized intensity in this region is reported as , whereas the significance of the JCMT polarized intensity in the head-tail structure in Sgr C is (Lu et al., 2024). is equivalent to a 10∘ uncertainty in the polarization angle, and the JCMT orientations disagree with those obtained from FIREPLACE by 70-90∘ in this Southern region.
To the Southwest of the head-tail structure our SOFIA/HAWC+ observations reveal a curved magnetic field structure at = 359.417∘, = -0.167∘. There are also JCMT pseudovectors at this location which are again highly spatially variable. The curvature in this region is similar to the field structures seen for the Brick and the TLP clouds that were attributed to out-flowing winds.A possible source of these winds is a compact source identified in the ATLASGAL survey (AGAL 359.474-0.152; Contreras et al., 2013), whose location is marked with a red ellipse in Figure 21. This source has an integrated flux of 168 Jy and a total gas mass of 5 104 M⊙ (Kendrew et al., 2013). These properties make AGAL359.474-0.152 favorable for forming massive stars (Lada et al., 2012).
Therefore, out-flowing winds from star formation in Sgr C, likely originating from AGAL 359.474-0.152 is a reasonable origin for the structure of the magnetic field in this region of Sgr C, especially in light of the ample evidence of star formation in Sgr C (e.g. Yusef-Zadeh et al., 2009; Lu et al., 2019). However, there are alternative mechanisms that could result in the magnetic field distribution seen in this region, such as shear from orbital motion of the structure.
As with the previously discussed clouds, the fractional polarizations derived from our 214 µm SOFIA/HAWC+ observations increase towards the periphery of Sgr C in lower emission regions. In the region of diffuse emission to the South of the head-tail structure fractional polarizations range from 4 – 5%. In the head-tail structure, conversely, the fractional polarizations are 1%, similar to what is observed in the peak emission regions of other prominent CMZ clouds.
@@ 6 CONCLUSIONS
In this work, the full FIREPLACE survey covering the entire CMZ polarimetrically at 214 µm is presented. First, the reduction method performed on the observations obtained in 2022 is detailed, comprising the second FIREPLACE data release (DR2) and observing the region of the CMZ extending from the Brick to Sgr C (corresponding to a 1∘ 0.75∘ region of the sky). The reduced DR2 observations are then combined with the DR1 observations presented in FIREPLACE I to obtain a polarimetric map of the entire CMZ at 214 µm (a 1.5∘ 0.75∘ region of the sky). CMZ-wide trends of the magnetic field are then inferred and analyzed, with the fields local to a set of prominent molecular clouds inspected in more detail. The following summarizes the results from the inspection of the full FIREPLACE survey:
@@
The sensitivity of our HAWC+ observations makes it possible to derive the polarization angle and magnetic field orientations in fainter molecular structures in the CMZ at a resolution of 19.6″ than ever before. The distribution of 64,000 Nyquist-sampled magnetic field pseudovectors from the full FIREPLACE survey is shown in Figure 10.
@@
A bimodal distribution in the CMZ-wide polarization angle orientations is identified from the full FIREPLACE survey. This bimodal distribution shows enhancements in directions parallel (approximately -16∘) and perpendicular (approximately -80∘) to the Galactic plane. This is a similar bimodal distribution to the one recovered in FIREPLACE I for the Sgr B2 region. The component that is parallel to the Galactic plane could either indicate a horizontal magnetic field component in the CMZ (Mangilli et al., 2019; Guan et al., 2021) or a component external to the CMZ. The bimodal component that is perpendicular to the Galactic plane could be related to the vertical magnetic field distribution in the GC (e.g. Morris, 2006), possibly revealing connections between the magnetic fields traced by infrared dust polarization and radio synchrotron emission.
Regions of vertical magnetic field (as shown in Figure 15) are not confined to the locations of detected NTFs, whereas regions where the magnetic field is parallel to the Galactic plane tend to be coincident with molecular clouds. This result supports the theory that there is a vertical magnetic field that pervades the GC and is sheared into a horizontal field in the higher density CMZ molecular clouds.
@@
The distributions of the magnetic fields of a set of prominent molecular clouds in the CMZ are also analyzed: the Brick, Three Little Pigs (TLP) cloud complex, 50 km s-1 cloud, CND, CO 0.02-0.02, 20 km s-1 cloud, and Sgr C. The magnetic field derived from FIREPLACE generally traces the morphology of the molecular clouds. In addition, there is evidence that expanding shock fronts from SNRs, advancing ionization fronts from H II regions, and tidal forces from orbital motion might cause compression and shearing leading to the organization of the magnetic fields in the Brick, TLP cloud complex, 50 km s-1 cloud, and Sgr C. However, not all of the prominent clouds studied in this work exhibit signatures of compression or winds (the 20 km s-1 cloud, for example). These results nonetheless indicate that the compression from winds and shear from tidal forces are important mechanisms for ordering the magnetic field distributions within CMZ clouds.
@@
The FIREPLACE observations presented in this work have the potential to enhance our understanding of the role magnetic fields play in the star formation process in the complex CMZ. Furthermore, the magnetic field distribution derived from these observations will improve our ability to develop a unified picture of the CMZ-wide magnetic field throughout the GC region. The reduced observations presented in this paper are publicly available on the NASA IPAC Legacy Surveys webpage as FITS files 555https://irsa.ipac.caltech.edu/data/SOFIA/docs/data/legacy-programs/two-color-polarimetric-survey-galactic-center-pilot-legacy-program/index.html.
Appendix A.1 Extended Polarization Artifacts in Different DRP Versions
We show an example of the polarization artifacts occurring in DRP version 3.2.0 but not in DRP version 2.7.0 for our FIREPLACE observations in Figure A.1. These artifacts appear in DRP version 2.7.0 when using the “-extended” flag, but do not appear without this flag666This difference could be a result of the updated image reconstruction SOFIA developed for version 3.2.0. The details of this update are outside of the scope of this paper but are discussed in the HAWC+ data reduction manuals: https://irsa.ipac.caltech.edu/data/SOFIA/docs/data/data-pipelines/. The FIREPLACE field shown in Figure A.1 was reduced using identical reduction parameters (as indicated in the caption of Figure A.1), with the only difference being the DRP version used to perform the reduction.
@@ Appendix A.2 Guide to Individual Fields and CAL Files
We tabulate the fields and calibrated files used for the DR2 FIREPLACE data set in Table A.1. The files in Table A.1 contain calibrated data with pointing corrections applied. Field numbers refer to the locations in Figure 2. “CCW” and “CW” refer to the rotation direction of the major scan direction relative to perpendicular to Galactic North (30∘ as discussed in Section 2.1). For completeness, files from Flight 890 are included in this table, though as discussed in the main paper these files are not used in the DR2 reduction. Horizontal lines are included to separate observations from different flights.
Table A.2 tabulates the FIREPLACE DR1 files that are combined with the DR2 files shown in Table A.1 to produce the full FIREPLACE data release covering the entire CMZ. The format of Table A.2 is identical to that of Table A.1 except that orientation (“CCW” or “CW”) is not specified since the orientation of the DR1 observations was more variant, as discussed in Section 2.1.
@@ Appendix A.3 Histograms of the 2022 Polarization Pseudovectors
To do so we sort the Nyquist-sampled polarization angles shown in Figure 3 into 10∘ bins. The histogram representation of the binned polarization angles for the DR2 observations is shown in Figure A.2. As with the histogram of the full FIREPLACE release, an angle of 0∘ indicates a magnetic field that is aligned with the Galactic plane, with an angle of 90∘ indicating a magnetic field that is oriented perpendicular to the Galactic plane.